OpenStudy (anonymous):

5.If the cumulative distribution function of X is given by the function below, then find P (X < 1.5). F(x)=0, if x ≤ 0 F(x)=1/8 x^3, if 0< x ≤ 2 F(x)=1, if x > 2

OpenStudy (anonymous):

Please help, I am a newbie with CDF and PDF. I am stuck with this question.

OpenStudy (anonymous):

Use these tools: http://www.acalculator.com/math-calculators.html

OpenStudy (anonymous):

Thanks, Let me see if the site can help.

OpenStudy (anonymous):

Nope, the site doesn't have tools needed to solve my problem.

OpenStudy (anonymous):

The CDF of a probability distribution is defined as \(F_X(x)=P(X\le k)\), so the probability you want is exactly \(F_X(1.5)\). Note: For continuous random variables, \(P(X\le k)=P(X<k)\).